Three point charges are located at the corners of a triangle: q1 =-5 μC at (-1,1), q2 = +3pC at ( Q3 = +2 μC at (1,1), with the coordinates given in m. Calculate the electric field E at (0,0) and the force F that would be exerted on a charge of q =-3pC at (0,0). 1-1), and

can anyone please explain how to solve this with steps?
Answer #1

Ea ante -SMC (1,1) NE A +2MC Buc +34ch Ez 100,0) E = K. 5x10-6 (52) E = 9x109 x 5x106 Es = 22:00 NIC En = 9x109x 3XJ0-6 10 =